### v2

# A Unified Instance Segmentation Framework for Completely Occluded Objects and Dense Objects in Robot Vision Measurement

Abstract
Instance segmentation for completely occluded objects and dense objects in robot vision measurement are two challenging tasks. To uniformly deal with them, this paper proposes a unified coarse-to-fine instance segmentation framework, CFNet, which uses box prompt-based segmentation foundation models (BSMs), e.g., Segment Anything Model. Specifically, CFNet first detects oriented bounding boxes (OBBs) to distinguish instances and provide coarse localization information. Then, it predicts OBB prompt-related masks for fine segmentation. CFNet performs instance segmentation with OBBs that only contain partial object boundaries on occluders to predict occluded object instances, which overcomes the difficulty of existing amodal instance segmentation methods in directly predicting occluded objects. In addition, since OBBs only serve as prompts, CFNet alleviates the over-dependence on bounding box detection performance of current instance segmentation methods using OBBs for dense objects. Moreover, to enable BSMs to handle OBB prompts, we propose a novel OBB prompt encoder. To make CFNet more lightweight, we perform knowledge distillation on it and introduce a Gaussian label smoothing method for teacher model outputs. Experiments demonstrate that CFNet outperforms current instance segmentation methods on both industrial and public datasets. The code is available at https://github.com/zhen6618/OBBInstanceSegmentation.
I INTRODUCTION
Instance segmentation provides foundational information for numerous tasks based on robot vision measurement, such as robot grasping [1, 2] and autonomous driving [3]. This work focuses on two challenging difficulties in robot vision measurement: completely occluded object instance segmentation and dense object instance segmentation. We aim to develop a unified framework to handle these difficulties.
Figure 1: Examples of completely occluded objects (a) and dense objects (b). Original objects are inside the corresponding black dotted boxes. (a): reference holes (blue) are occluded by bolts or nuts (i.e., occluders, shown in red). Oriented bounding boxes (green) contain occluder boundaries (orange) that are located at the contact surface between occluders and occluded reference holes. (b): dense vehicles (red) are surrounded by horizontal bounding boxes (yellow) or oriented bounding boxes (green).
Instance segmentation for completely occluded objects (called occludees) is difficult, since the information available for inferring occluded objects is very limited. For example, in Fig. 1(a), instance segmentation of reference holes that provide optimal assembly positions is usually required in robot assembly tasks. In several scenarios, reference holes (i.e., occludees) are screwed with bolts and nuts, completely occluded by these industrial parts (called occluders). To perform instance segmentation on occludees, current amodal instance segmentation methods [4, 5, 6, 7] directly predict occludees. Although they can automatically extract features related to occludees for inference, the performance is poor.
Dense object instance segmentation is a long-standing challenge [8] in robot vision measurement. For instance, in Fig. 1(b), objects are densely packed in multiple orientations in unmanned aerial vehicle (UAV) measurement. Since horizontal bounding boxes (HBBs) introduce many interference areas, oriented bounding boxes (OBBs) are often used for instance identification [9, 10, 11, 12], e.g., ISOP [12] predicts masks in the detected OBBs. However, bounding box-based instance segmentation methods are overly dependent on bounding box detection performance, especially for OBBs that are sensitive to both position and orientation. This dependence makes instance segmentation methods using OBBs more difficult and unstable.
In this paper, we propose CFNet, a unified coarse-to-fine framework for both completely occluded and dense object instance segmentation, which uses box prompt-based segmentation foundation models (BSMs), e.g., Segment Anything Model (SAM) [13]. CFNet first detects OBBs to distinguish instances, identify classes, and provide coarse localization information. Then, it predicts OBB-related segmentation masks. Our motivation is inspired by the powerful segmentation capabilities of BSMs and their excellent box prompt mechanisms. CFNet uses an OBB detection module (e.g., Oriented R-CNN [14]) for coarse detection, and leverages BSMs to predict OBB-related masks for fine segmentation.
CFNet improves BSMs in two aspects. First, to handle OBB prompts, CFNet introduces a novel OBB prompt encoder. Second, since BSMs usually have high computational complexity, we perform knowledge distillation on the box prompt encoders and mask decoders. Specifically, a Gaussian smoothing method for teacher model outputs is introduced. In this work, CFNet chooses SAM as the baseline model and makes improvements in the above two aspects.
Compared with amodal instance segmentation methods, CFNet avoids directly predicting occludee instances by performing instance segmentation with OBBs that only contain partial object boundaries on occluders to infer occludees. In addition, OBBs used in CFNet provide more accurate location information than HBBs used in amodal instance segmentation methods. Compared with current instance segmentation methods using OBBs for dense objects, CFNet only uses OBBs as prompts to guide object segmentation, so the segmentation results are less dependent on bounding box detection performance, which improves instance segmentation performance. Moreover, CFNet is based on BSMs that are pretrained on large-scale data for segmentation tasks, helping to enhance feature extraction and segmentation capabilities. The main contributions are summarized as follows.
  • We propose CFNet, a unified coarse-to-fine framework for both completely occluded and dense object instance segmentation in robot vision measurement. CFNet uses BSMs and outperforms current instance segmentation methods on both industrial and public datasets.
  • A novel OBB prompt encoder is proposed to effectively encode OBB prompts and guide BSMs to generate OBB prompt-related segmentation masks.
  • We propose a knowledge distillation method for the OBB prompt encoders and mask decoders of BSMs, significantly reducing the computational complexity with a minor loss in accuracy. Specifically, a Gaussian smoothing method for teacher model outputs is proposed.
II Related Work
II-A Amodal Instance Segmentation
Amodal instance segmentation directly predicts completely occluded object instances. The first amodal instance segmentation method [7] predicted the amodal masks and corresponding bounding boxes in an iterative bounding box expansion manner. Zhu et al. [15] directly trained general instance segmentation methods on amodal instance segmentation datasets. Based on Mask R-CNN [16], Qi et al. [5] predicted occluded parts by adding a multi-task branch with multi-level coding. Similarly, based on Mask R-CNN, ORCNN [6] first predicted amodal masks through an amodal mask prediction branch and then combined a visible mask prediction branch to predict occluded parts. Different from the above methods that only considered occludees, BCNet [4] proposed a bilayer graph convolutional network to simultaneously consider interactions between occluders and occludees. Since the information available for inferring occludees is very limited, these methods have poor performance.
II-B Instance Segmentation with Oriented Bounding Boxes
Compared with instance segmentation with HBBs [16, 17], instance segmentation with OBBs that have less interference areas provides more accurate location information. Based on Mask R-CNN, a Region of Interest (RoI) learner was applied to HBB proposals to generate OBB proposals, followed by a head that generated segmentation masks [12]. Based on a similar idea, Feng et al. [11] proposed a part-aware instance segmentation network with OBB proposals for bin picking. Follmann and König [10] first generated the final OBBs using a two-stage object detection approach, and then predicted masks within the detected OBBs. Rotated Blend Mask R-CNN [9] proposed a top-down and bottom-up structure for oriented instance segmentation. However, these methods are overly dependent on OBB detection performance, making predictions more difficult and unstable.
II-C Box Prompt-Based Segmentation Foundation Models
Benefiting from pretraining on large-scale data for segmentation tasks, BSMs exhibit powerful segmentation and generalization capabilities. For example, SAM [13] was trained on more than 1B segmentation masks from 11M images, showing remarkable segmentation and zero-shot generalization capabilities. A BSM for medical image segmentation is proposed in [18]. To enhance interactive performance, BSMs incorporate prompt mechanisms. A representative work is SAM, which took HBBs as prompts and predicted segmentation masks with respect to HBBs. MobileSAM [19] retained the HBB prompt mechanism and proposed a lightweight version of SAM for edge devices.
Figure 2: Architecture of the proposed CFNet. Compared with dense object instance segmentation, completely occluded object instance segmentation needs more post-processing steps to transform occluder instances into occludee instances.
II-D Box Prompt Encoder
Existing methods usually use HBBs as box prompts for BSMs, such as SAM and MobileSAM. The current designs of HBB prompt encoders are mainly inspired by SAM. A HBB was represented by two points, i.e., the top-left corner point and the bottom-right corner point. Then, a point was encoded as the sum of a Gaussian positional encoding [20] of the location and a learned embedding representing the top-left corner or bottom-right corner. Hence, a box was encoded as a combination of the two encoded point embeddings. However, to the best of our knowledge, there is no relevant research on OBB prompt encoders yet.
II-E Label Smoothing for Box Prompt encoders and Segmentation networks in Knowledge Distillation
In knowledge distillation for segmentation tasks, label smoothing helps reduce overfitting and enhances model performance. Inception-v3 [21] proposed uniform label smoothing (ULS), i.e., the weighted average of one-hot labels and uniform distributions. Spatially varying label smoothing (SVLS) [22] applied a discrete spatial Gaussian kernel to segmentation labels to smooth one-hot labels. These two label smoothing methods were commonly used in segmentation tasks. For example, Park et al. [23] applied ULS and proposed pixel-wise adaptive label smoothing for self-distillation. P-CD [24] utilized SVLS to generate segmentation boundary uncertainty and soft targets. Unlike applying SVLS to ground truth (GT) labels, we apply Gaussian smoothing to the segmentation masks of the teacher model, which yields better distillation performance. As for label smoothing methods used for box prompt encoders, to our knowledge, no relevant research has been published yet.
III Methods
III-A Overview
The structure of CFNet is shown in Fig. 2. It is mainly composed of four parts: an OBB detection module, an image encoder, an OBB prompt encoder, and a mask decoder. Specifically, images containing dense or occluded objects serve as input. The OBB detection module is used to detect OBBs that distinguish instances, identify classes, and provide coarse localization information. The image encoder is used to transform input images into high-dimensional feature representations and generate image embeddings. The OBB prompt encoder is responsible for encoding OBB prompts and generating prompt embeddings. Then, the image embeddings and prompt embeddings are combined in the mask decoder to predict OBB prompt-related segmentation masks.
For dense object instance segmentation, CFNet detects OBBs that encompass entire instances and predicts segmentation masks of dense objects. For completely occluded object instance segmentation, CFNet avoids directly predicting occludee instances by performing instance segmentation with OBBs that only contain partial object boundaries on occluders. Specifically, OBBs contain occluder boundaries that are located at the contact surface between occluders and occludees (see Fig. 1(a)). Then, post-processing steps transform occluder instances into partial boundary instances and use prior geometric relationships between boundary instances and occludee instances to obtain occludee instances (e.g., in robot assembly, the boundary shape of the occluded reference holes is an ellipse). However, partial bounding boxes are not suitable for current instance segmentation methods since their bounding boxes are usually designed to encompass entire instances.
Furthermore, knowledge distillation is applied to the OBB prompt encoder and mask decoder to reduce the computational complexity and make CFNet more lightweight. For OBB detection, we directly employ Oriented R-CNN [14], which is a two-stage OBB detection model with competitive detection accuracy and inference speed.
III-B Oriented Bounding Box Prompt Encoder
The flowchart of the proposed OBB prompt encoder is described in Fig. 3. The detected OBBs are generated by the OBB detection module. An OBB is first parameterized as ($x_{1},y_{1},x_{2},y_{2},\sin\theta,\cos\theta$), where ($x_{1},y_{1}$), ($x_{2},y_{2}$) and ($\sin\theta,\cos\theta$) represent the top-left corner point, bottom-right corner point and orientation, respectively. We further denote
$\bm{\varphi}_{1}=(x_{1},y_{1}),~{}~{}\bm{\varphi}_{2}=(x_{2},y_{2}),~{}~{}\bm{\theta}=(\sin\theta,\cos\theta),$
$\bm{\varphi}_{1}\in{[0,1)}^{2},~{}~{}\bm{\varphi}_{2}\in{[0,1)}^{2},~{}~{}\bm{\theta}\in{[0,1)}^{2},$ (1)
where ($\bm{\varphi}_{1},\bm{\varphi}_{2}$) and $\bm{\theta}$ are normalized pixel coordinates and normalized orientation coordinates, respectively. Considering the difference between position information ($\bm{\varphi}_{1},\bm{\varphi}_{2}$) and orientation information ($\bm{\theta}$), we first encode them separately and then fuse their encoded features.
Coordinate-based multilayer perceptrons (MLPs) take low-dimensional coordinates as input, such as $\bm{\varphi}_{1}$, $\bm{\varphi}_{2}$ and $\bm{\theta}$, and they have difficulty learning high-frequency information [20]. In CFNet, there are many MLP structures and transformer structures containing MLPs, so such a problem will degrade instance segmentation performance. To alleviate this problem, inspired by Gaussian positional encoding (GPE) for coordinates [20], CFNet introduces adaptive GPE (AGPE) to map the coordinate set ($\bm{\varphi}_{1},\bm{\varphi}_{2},\bm{\theta}$) into random Fourier features to better learn high-frequency information and enhance the performance of coordinate-based MLPs.
Given a d-dimensional coordinate $\bm{\tau}\in{[0,1)}^{d}$, the dense Fourier feature mapping is
$\gamma(\bm{\tau})=[$ $a_{0}\cos(2\pi\bm{G}_{0}^{T}\bm{\tau}),b_{0}\sin(2\pi\bm{G}_{0}^{T}\bm{\tau}),\ldots$
$a_{i}\cos(2\pi\bm{G}_{i}^{T}\bm{\tau}),b_{i}\sin(2\pi\bm{G}_{i}^{T}\bm{\tau}),\ldots]^{T},$ (2)
where $a_{i}$ and $b_{i}$ ($i=0,\ldots,\infty$) are Fourier series coefficients, and $\bm{G}_{i}\in\mathbb{R}^{d}$ is the corresponding Fourier basis frequency. Since random sparse sampling of Fourier features through an MLP matches the performance of using a dense sampling of Fourier features with the same MLP [20], AGPE sets all $a_{i}$ and $b_{i}$ to 1. $\bm{G}_{i}$ is sampled from a Gaussian distribution. Hence, the mapped random Fourier features are
$\gamma(\bm{\varphi}_{1})$ $=[\cos(2\pi G_{\bm{\varphi}_{1}}\bm{\varphi}_{1}),\sin(2\pi G_{\bm{\varphi}_{1}}\bm{\varphi}_{1})]^{T},$ (3)
$\gamma(\bm{\varphi}_{2})$ $=[\cos(2\pi G_{\bm{\varphi}_{2}}\bm{\varphi}_{2}),\sin(2\pi G_{\bm{\varphi}_{2}}\bm{\varphi}_{2})]^{T},$ (4)
$\gamma(\bm{\theta})$ $=[\cos(2\pi G_{\bm{\theta}}\bm{\theta}),\sin(2\pi G_{\bm{\theta}}\bm{\theta})]^{T},$ (5)
where
$G_{\bm{\varphi}_{1}}$ $\in\mathbb{R}^{(l_{p},2)}\sim\mathcal{N}(\bm{\rho}{\bm{\varphi}_{1}},\bm{\Sigma}_{\bm{\varphi}_{1}}),~{}~{}\gamma(\bm{\varphi}_{1})\in\mathbb{R}^{2l_{p}},$ (6)
$G_{\bm{\varphi}_{2}}$ $\in\mathbb{R}^{(l_{p},2)}\sim\mathcal{N}(\bm{\rho}{\bm{\varphi}_{2}},\bm{\Sigma}_{\bm{\varphi}_{2}}),~{}~{}\gamma(\bm{\varphi}_{2})\in\mathbb{R}^{2l_{p}},$ (7)
$G_{\bm{\theta}}$ $\in\mathbb{R}^{(l_{\theta},2)}\sim\mathcal{N}(\bm{\rho}{\bm{\theta}},\bm{\Sigma}_{\bm{\theta}})~{}~{}\gamma(\bm{\theta})\in\mathbb{R}^{2l_{\theta}}.$ (8)
The $l_{p}$ and $l_{\theta}$ are the lengths of the sampled Fourier basis frequencies. The $(\bm{\rho}{\bm{\varphi}_{1}},\bm{\rho}{\bm{\varphi}_{2}},\bm{\rho}{\bm{\theta}})$ and ($\bm{\bm{\Sigma}}_{\bm{\varphi}_{1}},\bm{\bm{\Sigma}}_{\bm{\varphi}_{2}},\bm{\bm{\Sigma}}_{\bm{\theta}}$) represent mean vectors and covariance matrices, respectively. To represent position information ($\gamma(\bm{\varphi}_{1})$, $\gamma(\bm{\varphi}_{2})$) and orientation information ($\gamma(\bm{\theta})$) in a unified representation space, we set $l_{p}=l_{\theta}$.
Figure 3: Architecture of the proposed OBB prompt encoder. The input is an OBB ($x,y,w,h,\theta$), where $(x,y)$, $w$, $h$ and $\theta$ represent the center point, width, height and orientation, respectively.
In many HBB prompt encoders of BSMs, $(\bm{\rho}{\bm{\varphi}_{1}},\bm{\rho}{\bm{\varphi}_{2}},\bm{\rho}{\bm{\theta}})$ and ($\bm{\bm{\Sigma}}_{\bm{\varphi}_{1}},\bm{\bm{\Sigma}}_{\bm{\varphi}_{2}},\bm{\bm{\Sigma}}_{\bm{\theta}}$) are considered as hyperparameters, which leads to difficulty in adapting to changes in datasets and searching optimal parameter settings. In CFNet, we adaptively learn ($\bm{\bm{\Sigma}}_{\bm{\varphi}_{1}},\bm{\bm{\Sigma}}_{\bm{\varphi}_{2}},\bm{\bm{\Sigma}}_{\bm{\theta}}$). $(\bm{\rho}{\bm{\varphi}_{1}},\bm{\rho}{\bm{\varphi}_{2}},\bm{\rho}{\bm{\theta}})$ are fixedly set to 0. Concretely, ($\bm{\bm{\Sigma}}_{\bm{\varphi}_{1}},\bm{\bm{\Sigma}}_{\bm{\varphi}_{2}},\bm{\bm{\Sigma}}_{\bm{\theta}}$) are the weights of a learnable embedding layer.
Inspired by SAM, the mapped random Fourier features ($\gamma(\bm{\varphi}_{1})$ and $\gamma(\bm{\varphi}_{2})$) are summed with two learned embeddings ($\bm{\omega}_{1}$ and $\bm{\omega}_{2}$) that represent the top-left corner point and bottom-right corner point in CFNet, respectively.
$\mathcal{E}(\bm{\varphi}_{1})=\gamma(\bm{\varphi}_{1})+\bm{\omega}_{1},$ (9)
$\mathcal{E}(\bm{\varphi}_{2})=\gamma(\bm{\varphi}_{2})+\bm{\omega}_{2}.$ (10)
Then, we obtain the encoded feature embeddings $\mathcal{E}(\bm{\varphi}_{p})$ with respect to position information, which is given as
$\mathcal{E}(\bm{\varphi}_{p})=\mathrm{Concat}[\mathcal{E}(\bm{\varphi}_{1}),~{}~{}\mathcal{E}(\bm{\varphi}_{2})].$ (11)
On the other hand, the parameterized orientation representation ($\sin\theta,\cos\theta$) of an OBB suffers from the boundary discontinuity problem [25]. Due to the periodicity of orientations, at boundary orientations (such as $0^{\circ}$ and $180^{\circ}$), small changes in orientations will result in large jumps in ($\sin\theta,\cos\theta$). To alleviate the boundary discontinuity problem, CFNet introduces learnable orientation correction embeddings $\bm{\omega}_{\theta}$. At the orientations where the boundary discontinuity problem occurs, $\bm{\omega}_{\theta}$ adaptively adjusts the effect of this problem. Hence, the encoded feature embeddings $\mathcal{E}(\bm{\theta})$ with respect to orientation information are
$\mathcal{E}(\bm{\theta})=\gamma(\bm{\theta})+\bm{\omega}_{\theta}.$ (12)
Finally, the encoded OBB prompt embeddings $\mathcal{P}_{\mathrm{OBB}}$ (also called prompt embeddings) are the concatenation of the encoded position feature embeddings and encoded orientation feature embeddings, which is described as
$\mathcal{P}_{\mathrm{OBB}}=\mathrm{Concat}[\mathcal{E}(\bm{\varphi}_{p}),~{}~{}\mathcal{E}(\bm{\theta})].$ (13)
Figure 4: The process of knowledge distillation for the OBB prompt encoder and mask decoder. “TE”, “BE” and “OE” represent encoded feature embeddings with respect to the top-left point, bottom-right point and orientation of an OBB, respectively. “GS” stands for Gaussian smoothing.
III-C Knowledge Distillation
BSMs exhibit powerful performance, however they usually have high computational complexity. Some methods, such as MobileSAM [19], focus on distilling image encoders. However, the performance gap between the distilled model and the teacher model is still large. To tackle this, CFNet uses the distilled image encoders, and further distills the prompt encoders and mask decoders of BSMs. Considering the powerful data-fitting capability of the teacher model, the student model only imitates the teacher model and does not learn from GT labels. Furthermore, to further enhance the performance and generalization capability of the student model, Gaussian smoothing is applied to teacher model outputs. The process of knowledge distillation for the proposed OBB prompt encoder and mask decoder is shown in Fig. 4.
We define a one-dimensional Gaussian kernel $\mathbf{G}^{1}_{k}(x_{i};\sigma)$ with a length of $k$, which is given as
$\mathbf{G}^{1}_{k}(x_{i};\sigma)=\frac{1}{\mathbf{G}^{1}_{s}}e^{-[\frac{(x_{i}-x_{c})^{2}}{2\sigma^{2}}]},$ (14)
where
$\mathbf{G}^{1}_{s}=\sum_{i}e^{-[\frac{(x_{i}-x_{c})^{2}}{2\sigma^{2}}]},~{}~{}i\in\{0,1,\ldots,k-1\}.$ (15)
The $x_{i}$ and $x_{c}$ represent the position of each element and the central position, respectively. $\sigma$ is the standard deviation. Given the encoded feature embeddings of the teacher model with respect to the top-left point ($\mathcal{E}^{t}(\bm{\varphi}_{1})\in\mathbb{R}^{2l}$, $l=l_{p}=l_{\theta}$), bottom-right point ($\mathcal{E}^{t}(\bm{\varphi}_{2})\in\mathbb{R}^{2l}$) and orientation ($\mathcal{E}^{t}(\bm{\theta})\in\mathbb{R}^{2l}$), we perform Gaussian smoothing by convolving the Gaussian kernel $\mathbf{G}^{1}_{k}$ with these encoded embeddings.
$\bm{G}^{t}(\bm{\varphi}_{1})$ $=\mathbf{G}^{1}_{k}\circledast\mathcal{E}^{t}(\bm{\varphi}_{1}),$ (16)
$\bm{G}^{t}(\bm{\varphi}_{2})$ $=\mathbf{G}^{1}_{k}\circledast\mathcal{E}^{t}(\bm{\varphi}_{2}),$ (17)
$\bm{G}^{t}(\bm{\theta})$ $=\mathbf{G}^{1}_{k}\circledast\mathcal{E}^{t}(\bm{\theta}),$ (18)
[…]
Then, the corresponding encoded feature embeddings ($\mathcal{E}^{s}(\bm{\varphi}_{1}),\mathcal{E}^{s}(\bm{\varphi}_{2}),\mathcal{E}^{s}(\bm{\theta})$) of the student model mimic the smoothed feature embeddings ($\bm{G}^{t}(\bm{\varphi}_{1}),\bm{G}^{t}(\bm{\varphi}_{2}),\bm{G}^{t}(\bm{\theta})$) of the teacher model. The mean squared error (MSE) loss ($\mathcal{L}_{\mathrm{prompt}}$) is used as the loss function.
$\mathcal{L}_{\mathrm{prompt}}$ $=\frac{1}{2l}{||\bm{G}^{t}(\bm{\varphi}_{1})-\mathcal{E}^{s}(\bm{\varphi}_{1})||}^{2}_{2}$
$+\frac{1}{2l}{||\bm{G}^{t}(\bm{\varphi}_{2})-\mathcal{E}^{s}(\bm{\varphi}_{2})||}^{2}_{2}$
$+\frac{1}{2l}{||\bm{G}^{t}(\bm{\theta})-\mathcal{E}^{s}(\bm{\theta})||}^{2}_{2}.$ (19)
For knowledge distillation on the mask decoder of a BSM, Gaussian smoothing makes the transition between boundary regions of different classes smoother and improves generalization capability. As the mask is two-dimensional, a two-dimensional Gaussian kernel $\mathbf{G}^{2}_{k\times k}(u_{i},v_{j};\delta)$ with a size of $k\times k$ is given as
$\mathbf{G}^{2}_{k\times k}(u_{i},v_{j};\delta)=\frac{1}{\mathbf{G}^{2}_{s}}e^{-[\frac{(u_{i}-u_{c})^{2}+(v_{j}-v_{c})^{2}}{2\delta^{2}}]},$ (20)
where
$\mathbf{G}^{2}_{s}=\sum_{i}\sum_{j}e^{-[\frac{(u_{i}-u_{c})^{2}+(v_{j}-v_{c})^{2}}{2\delta^{2}}]},~{}~{}i,j\in\{0,1,\ldots,k-1\}.$ (21)
The ($u_{i},v_{j}$) and ($u_{c},v_{c}$) represent the position of each element and the central position, respectively. $\delta$ is the standard deviation. In knowledge distillation for segmentation tasks, unlike current methods [22, 24] that apply Gaussian smoothing to GT labels, we apply it to masks generated by the teacher model.
Since the detected OBBs distinguish instances and identify classes, the output mask is only responsible for distinguishing between foreground and background and has a shape of ($H,W$), where $H$ and $W$ represent the height and width of the input image, respectively. Given the output $\bm{M}^{t}$ of the mask decoder of the teacher model, we first use the sigmoid function $\sigma(\cdot)$ to generate a foreground probability map, and then convolve it with the Gaussian kernel $\mathbf{G}^{2}_{k\times k}$.
$\bm{G}(\bm{M}^{t})=\sigma(\bm{M}^{t})\circledast\mathbf{G}^{2}_{k\times k},$ (22)
Subsequently, the generated smoothed target $\bm{G}(\bm{M}^{t})$ serves as the supervision target for the corresponding probability map ($\sigma(\bm{M}^{s})$) of the mask decoder of the student model. We use the binary cross entropy (BCE) loss ($\mathcal{L}_{\mathrm{mask}}$) as the loss function, which is described as
$\mathcal{L}_{\mathrm{mask}}=-\frac{1}{HW}\sum_{m^{t},m^{s}}[m^{t}\log m^{s}+(1-m^{t})\log(1-m^{s})],$
$m^{t}\in\bm{G}(\bm{M}^{t}),m^{s}\in\sigma(\bm{M}^{s}).$ (23)
Hence, the final distillation loss ($\mathcal{L}_{\mathrm{total}}$) is
$\mathcal{L}_{\mathrm{total}}=\lambda\cdot\mathcal{L}_{\mathrm{prompt}}+(1-\lambda)\cdot\mathcal{L}_{\mathrm{mask}},$ (24)
where $\lambda$ is the trade-off factor between the distillation loss of the OBB prompt encoder and that of the mask decoder.
IV Experiments
IV-A Dataset
To verify the effectiveness of the proposed methods, we conduct completely occluded object instance segmentation and dense object instance segmentation experiments on self-collected industrial and public datasets, respectively.
Figure 5: Self-designed robotic system for completely occluded object instance segmentation in the industrial robot assembly environment of the large commercial aircraft C919.
IV-A1 Industrial Dataset for Completely Occluded Objects
The self-designed industrial robotic vision system for large commercial aircraft assembly is shown in Fig. 5. It mainly consists of a UR10 robot, a vision sensor and a computer. Specifically, the UR10 robot is used to carry the vision sensor for operations. The computer is responsible for receiving and processing the image data collected by the vision sensor, and it runs the algorithm for completely occluded object instance segmentation. Due to the limited working space, reference holes are usually observed from the side.
All the images in the dataset are collected by this system, including three types of occluded reference holes, i.e., reference holes with bolts, nuts and untreated bolts (see the left image of Fig. 1). To reflect complex and changeable environments in the assembly process, each type of reference hole is photographed with the camera at various perspectives, distances and lighting conditions. We obtain a total of 800 images, with approximately the same number of each type. These images are divided into a training set, a validation set, and a testing set in a ratio of $7\colon 1\colon 2$ in a uniform sampling manner. All images are cropped to $1024\times 1024$. Then, these images are first flipped horizontally and vertically to quadruple the number. Experimental GT labels are obtained from industrial standard workpieces. Furthermore, the Mosaic method that randomly mixes four images is used for data augmentation.
IV-A2 Public Dataset for Dense Objects
iSAID [8] is a large-scale and densely annotated instance segmentation dataset in aerial images. It has 655,451 object instances across 15 categories. The training set, validation set and testing set contain 1411, 458 and 937 images, respectively. All images are cropped into $1024\times 1024$ patches with a stride of 500. Since iSAID dataset does not provide the GT parameters of OBBs, the smallest OBBs of instance masks are used as regression targets. We only choose densely distributed large vehicles, small vehicles and ships as targets for dense object instance segmentation.
Figure 6: Some visualization results on the completely occluded object instance segmentation dataset. Bounding boxes of different colors and segmentation masks of different colors represent different object classes and different instances, respectively.
IV-B Implementation Details
The training settings for Oriented R-CNN [14] are the same as the original method. In the OBB prompt encoder, the length of the sampled Fourier basis frequencies is 128. For knowledge distillation, the teacher model (called CFNet*) is based on SAM with the powerful ViT (vision transformer)-H image encoder and the proposed OBB prompt encoder. The student model (i.e., CFNet) uses a lightweight pretrained ViT-Tiny image encoder and the proposed OBB prompt encoder. Their mask decoders and OBB detection modules are the same. During training, the weights of their image encoders are frozen, and the other parts are fine-tuned. The teacher model has the same training loss as SAM and has been trained before distillation. The trade-off factor $\lambda$ for the distillation loss is 0.1. More implementation details can be found in our open-source code. All experiments are conducted on an RTX 3080Ti GPU. For evaluation, we use the standard COCO metrics: AP (i.e., AP50-95), AP50 and AP75, where “AP” means average precision over IoU (Intersection over Union) threshold.
IV-C Completely Occluded Object Instance Segmentation
In completely occluded object instance segmentation experiments, we compare CFNet, SAM, and amodal methods. SAM uses the ViT-H to encode images and the HBB detection branch of Mask R-CNN to generate HBB prompts (the same below). Amodal methods directly predict occludee instances. CFNet and SAM perform the proposed instance segmentation method on occluders to obtain occludee instances. We also include SAM after fine-tuning on this dataset and the teacher model (CFNet*) used for distillation in comparison. For fair comparison, we replace the backbones of current amodal instance segmentation methods with the ViT-H image encoder used in BSMs (indicated by $\dagger$).
TABLE I: Comparison results on the completely occluded object instance segmentation dataset.
Method
AP
AP50
AP75
FPS
AmodalMask [15]
19.49
35.18
24.36
16.1
ASN [5]
40.88
61.24
45.93
18.6
ORCNN [6]
37.60
57.78
40.16
22.2
BCNet [4]
45.30
65.13
51.79
7.4
[…]
50.09
69.30
56.42
2.7
SAM [13]
26.15
42.74
30.28
2.8
Fine-tuned SAM
68.66
82.23
71.38
2.8
CFNet*
75.13 89.21 77.98
2.7
CFNet
73.46
86.48
76.20
20.4
As shown in Table I, the accuracy of CFNet and CFNet* is superior to other methods, especially for high-precision metrics. Additionally, CFNet exhibits competitive inference speed. Some visualization results are presented in Fig. 6. Compared with amodal instance segmentation methods, CFNet provides an effective solution for completely occluded object instance segmentation. In addition, compared with fine-tuned SAM using HBBs, the accuracy of CFNet and CFNet* using OBBs is significantly improved, which indicates the necessity of using OBBs. Moreover, our proposed knowledge distillation method significantly reduces the computational complexity with a minor loss in accuracy.
Figure 7: Some visualization results on the dense object instance segmentation dataset iSAID. Bounding boxes of different colors and segmentation masks of different colors represent different object classes and different instances, respectively.
IV-D Dense Object Instance Segmentation
For dense object instance segmentation, we compare CFNet, SAM (with HBBs), and instance segmentation methods using OBBs. We also include fine-tuned SAM and the teacher model used for distillation in comparison, and replace the backbones of current instance segmentation methods using OBBs with the ViT-H image encoder used in BSMs.
TABLE II: Comparison results on the dense object instance segmentation dataset. “RB Mask R-CNN” represents Rotated Blend Mask R-CNN [9].
Method
AP
AP50
AP75
FPS
OIS [10]
27.41
47.85
30.71
18.4
RB Mask R-CNN [9]
29.07
50.96
32.64
16.3
ISOP [12]
34.73
59.16
36.84
19.6
[…]
37.81
61.95
39.45
2.7
SAM [13]
14.37
41.88
18.52
2.8
Fine-tuned SAM
20.26
45.30
24.59
2.8
CFNet*
45.10 66.45 47.92
2.7
CFNet
43.76
64.29
46.13
20.4
As presented in Table II, CFNet and CFNet* outperform other methods, and CFNet achieves the highest inference speed of 20.4 FPS. In addition, CFNet achieves a remarkable dense object instance segmentation accuracy of 43.76% AP, 64.29% AP50, and 46.13% AP75. Fig. 7 shows some dense object instance segmentation results.
Compared with existing instance segmentation methods using OBBs, CFNet achieves higher accuracy since it reduces the dependence on OBBs and is based on BSMs. Significant accuracy improvements over fine-tuned SAM also indicate the benefit of using OBBs. Moreover, compared with CFNet*, the notable inference speed improvement and minor accuracy loss of CFNet also demonstrate the effectiveness of the proposed knowledge distillation method.
IV-E Ablation Studies
To evaluate the effectiveness of the components of CFNet, a series of ablation experiments are conducted on the iSAID dataset. We use the most commonly used AP50 as the metric.
IV-E1 Oriented Bounding Box Prompt Encoder
The OBB prompt encoder introduces positional encoding and specific embeddings to encode OBBs. In Table III, “OCE”, “TCE” and “BCE” represent the specific orientation correction embeddings, top-left corner embeddings and bottom-right corner embeddings, respectively. Compared with GPE, AGPE encodes location information better by adaptively learning parameters of Gaussian distributions. Specific embeddings further improve instance segmentation performance.
TABLE III: Performance of each component of the OBB prompt encoder.
GPE
AGPE
TCE+BCE
OCE
AP50
61.40
62.27
63.56
64.29
TABLE IV: Comparison of different learning targets and label smoothing methods in knowledge distillation.
GT+ULS
GT+GS
T+GT
T
T+ULS
T+GS
63.51
63.72
64.01
64.07
64.15
64.19
64.24
IV-E2 Knowledge Distillation
We first compare the performance of CFNet learning from the teacher model and GT labels, and test the impact of different label smoothing methods. “T” represents learning from the teacher model. The smoothing factor in ULS is 0.1. Gaussian smoothing (GS) has the same settings as SVLS, with a Gaussian kernel size of $3\times 3$ and a standard deviation of 1.0. The weight factor of the teacher model and GT labels is 0.1. “${}^{-}$” represents distillation only for the mask decoder.
The experimental results in Table IV demonstrate that learning from the teacher model has higher performance than learning from GT labels (learning only from GT labels means no knowledge distillation is performed), which is mainly thanks to the powerful data-fitting capabilities of BSMs. Furthermore, adding distillation for the OBB prompt encoder effectively enhances performance. Moreover, Gaussian smoothing further improves instance segmentation performance and is superior to ULS. In addition, we compare the impact of different parameter settings in Gaussian smoothing on instance segmentation performance in Table V. When $k=5$, $\sigma=0.3$, and $\delta=1.0$, the highest instance segmentation accuracy of 64.29% AP50 is achieved.
TABLE V: Impact of parameter settings in Gaussian smoothing.
$\sigma=0.3$ $\sigma=0.5$
$\delta=0.5$
$\delta=1.0$
$\delta=0.5$
$\delta=1.0$
$k=3$
64.24
64.24
64.20
64.17
$k=5$
64.22
64.29
64.18
64.21
$k=7$
64.19
64.25
64.15
64.14
IV-E3 Oriented Bounding Box Detection
To explore the impact of the OBB detection module on the instance segmentation performance, we compare Oriented R-CNN used in CFNet with some other OBB detection methods [26, 27]. In addition, a HBB detection method (HBB detection branch of Mask R-CNN) is also added for comparison. As shown in Table VI, the effective use of OBBs significantly improves instance segmentation accuracy, and CFNet is not sensitive to specific OBB detection methods.
TABLE VI: Performance comparison of OBB and HBB detectors.
Mask
Oriented
Oriented
S2ANet [27]
R-CNN [16]
R-CNN [14]
Reppoints [26]
42.95
64.29
64.37
64.06
V Conclusions
This paper proposes CFNet, a unified coarse-to-fine framework for both completely occluded and dense object instance segmentation in robot vision measurement. CFNet uses BSMs and introduces a novel OBB prompt encoder. To make CFNet more lightweight, a knowledge distillation method involving Gaussian smoothing for teacher model outputs is applied to the OBB prompt encoder and mask decoder. Experiments on both industrial and public datasets demonstrate that CFNet outperforms current instance segmentation methods.
One limitation of this work is that CFNet relies on prior geometric properties to achieve completely occluded object instance segmentation. In the future, we plan to utilize more types of prior knowledge to improve the robustness. It is worth noting that CFNet not only uniformly handles the two difficulties of completely occluded and dense object instance segmentation, but also can be used for general instance segmentation tasks. Moreover, CFNet can also handle objects that are simultaneously in completely occluded and dense environments, such as instance segmentation of containers with UAVs. We also plan to explore more applications of CFNet in various complex robotic tasks.

### v3

@@
-A Unified Instance Segmentation Framework for Completely Occluded Objects and Dense Objects in Robot Vision Measurement
+An Efficient Instance Segmentation Framework Based on Oriented Bounding Boxes
 Abstract
@@
 TABLE I: Comparison results on the completely occluded object instance segmentation dataset.
-Method
-AP
-AP50
-AP75
-FPS
-AmodalMask [15]
-19.49
-35.18
-24.36
-16.1
-ASN [5]
-40.88
-61.24
-45.93
-18.6
-ORCNN [6]
-37.60
-57.78
-40.16
-22.2
-BCNet [4]
-45.30
-65.13
-51.79
-7.4
+Method AP AP50 AP75 FPS
+AmodalMask [15] 19.49 35.18 24.36 16.1
+ASN [5] 40.88 61.24 45.93 18.6
+ORCNN [6] 37.60 57.78 40.16 22.2
+BCNet [4] 45.30 65.13 51.79 7.4
 […]
-50.09
-69.30
-56.42
-2.7
-SAM [13]
-26.15
-42.74
-30.28
-2.8
-Fine-tuned SAM
-68.66
-82.23
-71.38
-2.8
-CFNet*
-75.13 89.21 77.98
-2.7
-CFNet
-73.46
-86.48
-76.20
-20.4
+SAM [13] 26.15 42.74 30.28 2.8
+Fine-tuned SAM 68.66 82.23 71.38 2.8
+CFNet* 75.13 89.21 77.98 2.7
+CFNet 73.46 86.48 76.20 20.4
 As shown in Table I, the accuracy of CFNet and CFNet* is superior to other methods, especially for high-precision metrics.
@@
 TABLE II: Comparison results on the dense object instance segmentation dataset. “RB Mask R-CNN” represents Rotated Blend Mask R-CNN [9].
-Method
-AP
-AP50
-AP75
-FPS
-OIS [10]
-27.41
-47.85
-30.71
-18.4
-RB Mask R-CNN [9]
-29.07
-50.96
-32.64
-16.3
-ISOP [12]
-34.73
-59.16
-36.84
-19.6
+Method AP AP50 AP75 FPS
+OIS [10] 27.41 47.85 30.71 18.4
+RB Mask R-CNN [9] 29.07 50.96 32.64 16.3
+ISOP [12] 34.73 59.16 36.84 19.6
 […]
-37.81
-61.95
-39.45
-2.7
-SAM [13]
-14.37
-41.88
-18.52
-2.8
-Fine-tuned SAM
-20.26
-45.30
-24.59
-2.8
-CFNet*
-45.10 66.45 47.92
-2.7
-CFNet
-43.76
-64.29
-46.13
-20.4
+SAM [13] 14.37 41.88 18.52 2.8
+Fine-tuned SAM 20.26 45.30 24.59 2.8
+CFNet* 45.10 66.45 47.92 2.7
+CFNet 43.76 64.29 46.13 20.4
 As presented in Table II, CFNet and CFNet* outperform other methods, and CFNet achieves the highest inference speed of 20.4 FPS.
@@
 TABLE III: Performance of each component of the OBB prompt encoder.
-GPE
-AGPE
-TCE+BCE
-OCE
-AP50
+GPE AGPE TCE+BCE OCE AP50
 61.40
@@
 TABLE IV: Comparison of different learning targets and label smoothing methods in knowledge distillation.
-GT+ULS
-GT+GS
-T+GT
-T
-T+ULS
-T+GS
-63.51
-63.72
-64.01
-64.07
-64.15
-64.19
-64.24
+GT+ULS GT+GS T+GT T T+ULS T+GS- T+GS
+63.51 63.72 64.01 64.07 64.15 64.19 64.24
 IV-E2 Knowledge Distillation
@@
 The weight factor of the teacher model and GT labels is 0.1.
-“${}^{-}$” represents distillation only for the mask decoder.
+“GS-” represents distillation only for the mask decoder.
 The experimental results in Table IV demonstrate that learning from the teacher model has higher performance than learning from GT labels (learning only from GT labels means no knowledge distillation is performed), which is mainly thanks to the powerful data-fitting capabilities of BSMs.
@@
 $\sigma=0.3$ $\sigma=0.5$
-$\delta=0.5$
-$\delta=1.0$
-$\delta=0.5$
-$\delta=1.0$
-$k=3$
-64.24
-64.24
-64.20
-64.17
-$k=5$
-64.22
-64.29
-64.18
-64.21
-$k=7$
-64.19
-64.25
-64.15
-64.14
+$\delta=0.5$ $\delta=1.0$ $\delta=0.5$ $\delta=1.0$
+$k=3$ 64.24 64.24 64.20 64.17
+$k=5$ 64.22 64.29 64.18 64.21
+$k=7$ 64.19 64.25 64.15 64.14
 IV-E3 Oriented Bounding Box Detection
@@
 TABLE VI: Performance comparison of OBB and HBB detectors.
-Mask
-Oriented
-Oriented
-S2ANet [27]
-R-CNN [16]
-R-CNN [14]
-Reppoints [26]
-42.95
-64.29
-64.37
-64.06
+Mask Oriented Oriented S2ANet [27]
+R-CNN [16] R-CNN [14] Reppoints [26]
+42.95 64.29 64.37 64.06
 V Conclusions

### v6

@@
-An Efficient Instance Segmentation Framework Based on Oriented Bounding Boxes
+OBSeg: Accurate and Fast Instance Segmentation Framework Using Segmentation Foundation Models with Oriented Bounding Box Prompts
 Abstract
-Instance segmentation for completely occluded objects and dense objects in robot vision measurement are two challenging tasks.
-To uniformly deal with them, this paper proposes a unified coarse-to-fine instance segmentation framework, CFNet, which uses box prompt-based segmentation foundation models (BSMs), e.g., Segment Anything Model.
-Specifically, CFNet first detects oriented bounding boxes (OBBs) to distinguish instances and provide coarse localization information.
+Instance segmentation in remote sensing images is a long-standing challenge.
+Since horizontal bounding boxes introduce many interference objects, oriented bounding boxes (OBBs) are usually used for instance identification.
+However, based on “segmentation within bounding box” paradigm, current instance segmentation methods using OBBs are overly dependent on bounding box detection performance.
+To tackle this problem, this paper proposes OBSeg, an accurate and fast instance segmentation framework using OBBs.
+OBSeg is based on box prompt-based segmentation foundation models (BSMs), e.g., Segment Anything Model.
+Specifically, OBSeg first detects OBBs to distinguish instances and provide coarse localization information.
 Then, it predicts OBB prompt-related masks for fine segmentation.
-CFNet performs instance segmentation with OBBs that only contain partial object boundaries on occluders to predict occluded object instances, which overcomes the difficulty of existing amodal instance segmentation methods in directly predicting occluded objects.
-In addition, since OBBs only serve as prompts, CFNet alleviates the over-dependence on bounding box detection performance of current instance segmentation methods using OBBs for dense objects.
-Moreover, to enable BSMs to handle OBB prompts, we propose a novel OBB prompt encoder.
-To make CFNet more lightweight, we perform knowledge distillation on it and introduce a Gaussian label smoothing method for teacher model outputs.
-Experiments demonstrate that CFNet outperforms current instance segmentation methods on both industrial and public datasets.
+Since OBBs only serve as prompts, OBSeg alleviates the over-dependence on bounding box detection performance of current instance segmentation methods using OBBs.
+Thanks to OBB prompts, OBSeg outperforms other current BSM-based methods using HBBs.
+In addition, to enable BSMs to handle OBB prompts, we propose a novel OBB prompt encoder.
+To make OBSeg more lightweight and further improve the performance of lightweight distilled BSMs, a Gaussian smoothing-based knowledge distillation method is introduced.
+Experiments demonstrate that OBSeg outperforms current instance segmentation methods on multiple datasets in terms of instance segmentation accuracy and has competitive inference speed.
 The code is available at https://github.com/zhen6618/OBBInstanceSegmentation.
+Index Terms:
+Instance segmentation in remote sensing images, oriented bounding boxes, segmentation foundation models, knowledge distillation, oriented bounding box prompt encoder.
 I INTRODUCTION
-Instance segmentation provides foundational information for numerous tasks based on robot vision measurement, such as robot grasping [1, 2] and autonomous driving [3].
-This work focuses on two challenging difficulties in robot vision measurement: completely occluded object instance segmentation and dense object instance segmentation.
-We aim to develop a unified framework to handle these difficulties.
-Figure 1: Examples of completely occluded objects (a) and dense objects (b).
-Original objects are inside the corresponding black dotted boxes.
-(a): reference holes (blue) are occluded by bolts or nuts (i.e., occluders, shown in red).
-Oriented bounding boxes (green) contain occluder boundaries (orange) that are located at the contact surface between occluders and occluded reference holes.
-(b): dense vehicles (red) are surrounded by horizontal bounding boxes (yellow) or oriented bounding boxes (green).
-Instance segmentation for completely occluded objects (called occludees) is difficult, since the information available for inferring occluded objects is very limited.
-For example, in Fig.
-1(a), instance segmentation of reference holes that provide optimal assembly positions is usually required in robot assembly tasks.
-In several scenarios, reference holes (i.e., occludees) are screwed with bolts and nuts, completely occluded by these industrial parts (called occluders).
-To perform instance segmentation on occludees, current amodal instance segmentation methods [4, 5, 6, 7] directly predict occludees.
-Although they can automatically extract features related to occludees for inference, the performance is poor.
-Dense object instance segmentation is a long-standing challenge [8] in robot vision measurement.
-For instance, in Fig.
-1(b), objects are densely packed in multiple orientations in unmanned aerial vehicle (UAV) measurement.
-Since horizontal bounding boxes (HBBs) introduce many interference areas, oriented bounding boxes (OBBs) are often used for instance identification [9, 10, 11, 12], e.g., ISOP [12] predicts masks in the detected OBBs.
-However, bounding box-based instance segmentation methods are overly dependent on bounding box detection performance, especially for OBBs that are sensitive to both position and orientation.
-This dependence makes instance segmentation methods using OBBs more difficult and unstable.
-In this paper, we propose CFNet, a unified coarse-to-fine framework for both completely occluded and dense object instance segmentation, which uses box prompt-based segmentation foundation models (BSMs), e.g., Segment Anything Model (SAM) [13].
-CFNet first detects OBBs to distinguish instances, identify classes, and provide coarse localization information.
+Instance segmentation of remote sensing images provides pixel-level foundational information for numerous tasks, such as maritime security [1, 2], agricultural measurement [3], urban management, and disaster assessment.
+Pixel-level perception enables related instance segmentation methods to depict the exterior appearance and contour of the object.
+However, objects in remote sensing images usually have the characteristics of multi-scale, dense arrangement, and arbitrary orientations, which brings a long-standing challenge to the instance segmentation task.
+Since objects are densely packed in multiple orientations, which introduces many interference regions into horizontal bounding boxes (HBBs) [4, 5], oriented bounding boxes (OBBs) are usually used for instance identification [6, 7, 8, 9] (see Fig.
+1).
+Existing instance segmentation methods are mainly based on the “segmentation within bounding box” paradigm, that is, object segmentation is mainly performed in the detected bounding box (ignoring small differences caused by RoIAlign).
+This paradigm makes the segmentation performance overly dependent on the bounding box detection performance.
+However, the core of the instance segmentation task is to distinguish different object instances and segment each instance with a corresponding mask.
+Bounding boxes can be used for instance identification and roughly locating instances, so bounding boxes can be regarded as tools to assist instance segmentation.
+If the segmentation is mainly restricted to the inside of the bounding boxes, the performance of instance segmentation will be degraded.
+Furthermore, since OBBs are sensitive to both position and orientation and are difficult to detect, this dependence makes instance segmentation methods using OBBs more challenging and unstable.
+Figure 1: For instance segmentation in remote sensing images, (a): HBB introduces many interference objects.
+(b): The “segmentation within bounding box” paradigm limits the segmentation to be performed mainly within the detected OBB, making the segmentation performance overly dependent on the OBB detection performance.
+Once the OBB detection is inaccurate, the mask segmentation will also be affected.
+(c) The proposed OBSeg only uses OBB as a prompt to guide object segmentation, so the segmentation result is less dependent on OBB detection performance.
+Although the OBB detection is inaccurate, the mask can be segmented accurately.
+In this paper, to reduce dependence on bounding box detection performance, we propose OBSeg, an accurate and fast instance segmentation framework that uses box prompt-based segmentation foundation models (BSMs), e.g., Segment Anything Model (SAM) [10].
+Our motivation is inspired by the powerful segmentation capabilities of BSMs and their excellent box prompt mechanisms.
+Although there are currently some instance segmentation methods [11] using BSMs (e.g., RSPrompter [11] improves on SAM), they use HBBs as prompts, introducing many interference regions.
+Different from BSMs with HBB prompts, OBSeg utilizes more accurate OBBs as prompts, which exhibits better segmentation performance.
+Specifically, OBSeg first detects OBBs to distinguish instances, identify classes, and provide coarse localization information.
 Then, it predicts OBB-related segmentation masks.
-Our motivation is inspired by the powerful segmentation capabilities of BSMs and their excellent box prompt mechanisms.
-CFNet uses an OBB detection module (e.g., Oriented R-CNN [14]) for coarse detection, and leverages BSMs to predict OBB-related masks for fine segmentation.
-CFNet improves BSMs in two aspects.
-First, to handle OBB prompts, CFNet introduces a novel OBB prompt encoder.
-Second, since BSMs usually have high computational complexity, we perform knowledge distillation on the box prompt encoders and mask decoders.
-Specifically, a Gaussian smoothing method for teacher model outputs is introduced.
-In this work, CFNet chooses SAM as the baseline model and makes improvements in the above two aspects.
-Compared with amodal instance segmentation methods, CFNet avoids directly predicting occludee instances by performing instance segmentation with OBBs that only contain partial object boundaries on occluders to infer occludees.
-In addition, OBBs used in CFNet provide more accurate location information than HBBs used in amodal instance segmentation methods.
-Compared with current instance segmentation methods using OBBs for dense objects, CFNet only uses OBBs as prompts to guide object segmentation, so the segmentation results are less dependent on bounding box detection performance, which improves instance segmentation performance.
-Moreover, CFNet is based on BSMs that are pretrained on large-scale data for segmentation tasks, helping to enhance feature extraction and segmentation capabilities.
+OBSeg uses an OBB detection module [12, 13, 14] for coarse detection, and leverages BSMs to predict OBB-related masks for fine segmentation.
+OBSeg improves BSMs in two aspects.
+First, to handle OBB prompts, OBSeg introduces a novel OBB prompt encoder.
+Second, since BSMs usually have high computational complexity, we perform knowledge distillation [15] on BSMs to improve the practicality.
+Although there are many methods [16, 17, 18] that use knowledge distillation to introduce lightweight BSMs, the performance gap between the distilled model and the teacher model needs to be further reduced.
+In OBSeg, a Gaussian smoothing method for teacher model outputs is introduced to further improve distillation performance.
+Compared with the “segmentation within bounding box” paradigm, OBSeg only uses OBBs as prompts to guide object segmentation, so the segmentation results are less dependent on bounding box detection performance, which effectively improves instance segmentation performance.
+Moreover, after distillation, OBSeg is more lightweight.
 The main contributions are summarized as follows.
-  • We propose CFNet, a unified coarse-to-fine framework for both completely occluded and dense object instance segmentation in robot vision measurement.
-CFNet uses BSMs and outperforms current instance segmentation methods on both industrial and public datasets.
+  1. 1.
+    We propose OBSeg, an accurate and fast instance segmentation framework using BSMs with OBB prompts.
+OBSeg outperforms current instance segmentation methods on multiple datasets in terms of instance segmentation accuracy and has competitive inference speed.
+  2. 2.
     A novel OBB prompt encoder is proposed to effectively encode OBB prompts and guide BSMs to generate OBB prompt-related segmentation masks.
-  • We propose a knowledge distillation method for the OBB prompt encoders and mask decoders of BSMs, significantly reducing the computational complexity with a minor loss in accuracy.
-Specifically, a Gaussian smoothing method for teacher model outputs is proposed.
+  3. 3.
+    A Gaussian smoothing-based knowledge distillation method is introduced to improve the instance segmentation performance of lightweight distilled BSMs.
 II Related Work
-II-A Amodal Instance Segmentation
-Amodal instance segmentation directly predicts completely occluded object instances.
-The first amodal instance segmentation method [7] predicted the amodal masks and corresponding bounding boxes in an iterative bounding box expansion manner.
-Zhu et al.
-[15] directly trained general instance segmentation methods on amodal instance segmentation datasets.
-Based on Mask R-CNN [16], Qi et al.
-[5] predicted occluded parts by adding a multi-task branch with multi-level coding.
-Similarly, based on Mask R-CNN, ORCNN [6] first predicted amodal masks through an amodal mask prediction branch and then combined a visible mask prediction branch to predict occluded parts.
-Different from the above methods that only considered occludees, BCNet [4] proposed a bilayer graph convolutional network to simultaneously consider interactions between occluders and occludees.
-Since the information available for inferring occludees is very limited, these methods have poor performance.
-II-B Instance Segmentation with Oriented Bounding Boxes
-Compared with instance segmentation with HBBs [16, 17], instance segmentation with OBBs that have less interference areas provides more accurate location information.
-Based on Mask R-CNN, a Region of Interest (RoI) learner was applied to HBB proposals to generate OBB proposals, followed by a head that generated segmentation masks [12].
-Based on a similar idea, Feng et al.
-[11] proposed a part-aware instance segmentation network with OBB proposals for bin picking.
+II-A Instance Segmentation with Oriented Bounding Boxes
+Compared with instance segmentation methods using HBBs [19, 20], instance segmentation methods using OBBs that have fewer interference regions provide more accurate location information.
+Most of existing methods build upon the “segmentation within bounding box” paradigm and the improved Mask R-CNN [19] framework.
+For example, a Region of Interest (RoI) learner was applied to HBB proposals to generate OBB proposals in ISOP [9], followed by a head that generated segmentation masks for OBB proposals.
 Follmann and König [7] first generated the final OBBs using a two-stage object detection approach, and then predicted masks within the detected OBBs.
 Rotated Blend Mask R-CNN [6] proposed a top-down and bottom-up structure for oriented instance segmentation.
-However, these methods are overly dependent on OBB detection performance, making predictions more difficult and unstable.
-II-C Box Prompt-Based Segmentation Foundation Models
+Although the above methods effectively improved current instance segmentation performance in remote sensing images, they are overly dependent on OBB detection performance, making predictions more difficult and unstable.
+Our proposed OBSeg uses OBBs as prompts to guide object segmentation, which effectively reduces this dependence.
+II-B Box Prompt-Based Segmentation Foundation Models
 Benefiting from pretraining on large-scale data for segmentation tasks, BSMs exhibit powerful segmentation and generalization capabilities.
@@
 A BSM for medical image segmentation is proposed in [21].
-To enhance interactive performance, BSMs incorporate prompt mechanisms.
+To enhance interactive performance, BSMs introduce prompt mechanisms.
 A representative work is SAM, which took HBBs as prompts and predicted segmentation masks with respect to HBBs.
-MobileSAM [19] retained the HBB prompt mechanism and proposed a lightweight version of SAM for edge devices.
-Figure 2: Architecture of the proposed CFNet.
-Compared with dense object instance segmentation, completely occluded object instance segmentation needs more post-processing steps to transform occluder instances into occludee instances.
-II-D Box Prompt Encoder
-Existing methods usually use HBBs as box prompts for BSMs, such as SAM and MobileSAM.
+To improve the practicality of BSMs, EfficientSAM [16] and MobileSAM [17] used knowledge distillation to introduce lightweight image encoders for edge devices.
+EdgeSAM [18] performs distillation on the image encoder, prompt encoder, and mask decoder.
+However, the performance of these distilled BSMs is limited.
+We introduce a Gaussian smoothing method for teacher model outputs to further improve distillation performance for BSMs.
+Figure 2: Architecture of the proposed OBSeg.
+It is mainly composed of four parts: an OBB detection module, an image encoder, an OBB prompt encoder, and a mask decoder.
+OBSeg first detects OBBs to distinguish instances, identify classes, and provide coarse localization information.
+Then, the mask decoder utilizes the image embeddings generated by the image encoder and the OBB prompt embeddings generated by the OBB prompt encoder to generate segmentation masks.
+In addition, Gaussian smoothing-based knowledge distillation is performed on the OBB prompt encoder and the mask decoder to make OBSeg more lightweight.
+II-C Box Prompt Encoder
+Existing methods usually use HBBs as box prompts for BSMs, such as SAM and EfficientSAM.
 The current designs of HBB prompt encoders are mainly inspired by SAM.
@@
 Hence, a box was encoded as a combination of the two encoded point embeddings.
-However, to the best of our knowledge, there is no relevant research on OBB prompt encoders yet.
-II-E Label Smoothing for Box Prompt encoders and Segmentation networks in Knowledge Distillation
+In remote sensing images, using OBBs helps improve instance segmentation performance.
+However, to the best of our knowledge, there is no relevant research on OBB prompt encoders yet, which is necessary for BSMs with OBB prompts.
+II-D Label Smoothing for Segmentation Networks in Knowledge Distillation
 In knowledge distillation for segmentation tasks, label smoothing helps reduce overfitting and enhances model performance.
@@
 P-CD [26] utilized SVLS to generate segmentation boundary uncertainty and soft targets.
-Unlike applying SVLS to ground truth (GT) labels, we apply Gaussian smoothing to the segmentation masks of the teacher model, which yields better distillation performance.
-As for label smoothing methods used for box prompt encoders, to our knowledge, no relevant research has been published yet.
+Different from applying SVLS to ground truth (GT) labels, to further improve the generalization capability and instance segmentation performance of the student model, we apply Gaussian smoothing to the segmentation masks of the teacher model, which yields better distillation performance.
 III Methods
 III-A Overview
-The structure of CFNet is shown in Fig.
+The structure of OBSeg is shown in Fig.
 2.
 It is mainly composed of four parts: an OBB detection module, an image encoder, an OBB prompt encoder, and a mask decoder.
-Specifically, images containing dense or occluded objects serve as input.
-The OBB detection module is used to detect OBBs that distinguish instances, identify classes, and provide coarse localization information.
+Specifically, the OBB detection module is used to detect OBBs that distinguish instances, identify classes, and provide coarse localization information.
 The image encoder is used to transform input images into high-dimensional feature representations and generate image embeddings.
@@
 Then, the image embeddings and prompt embeddings are combined in the mask decoder to predict OBB prompt-related segmentation masks.
-For dense object instance segmentation, CFNet detects OBBs that encompass entire instances and predicts segmentation masks of dense objects.
-For completely occluded object instance segmentation, CFNet avoids directly predicting occludee instances by performing instance segmentation with OBBs that only contain partial object boundaries on occluders.
-Specifically, OBBs contain occluder boundaries that are located at the contact surface between occluders and occludees (see Fig.
-1(a)).
-Then, post-processing steps transform occluder instances into partial boundary instances and use prior geometric relationships between boundary instances and occludee instances to obtain occludee instances (e.g., in robot assembly, the boundary shape of the occluded reference holes is an ellipse).
-However, partial bounding boxes are not suitable for current instance segmentation methods since their bounding boxes are usually designed to encompass entire instances.
-Furthermore, knowledge distillation is applied to the OBB prompt encoder and mask decoder to reduce the computational complexity and make CFNet more lightweight.
-For OBB detection, we directly employ Oriented R-CNN [14], which is a two-stage OBB detection model with competitive detection accuracy and inference speed.
+Furthermore, Gaussian smoothing-based knowledge distillation is applied to the OBB prompt encoder and mask decoder to make OBSeg more lightweight and practical.
 III-B Oriented Bounding Box Prompt Encoder
@@
 Coordinate-based multilayer perceptrons (MLPs) take low-dimensional coordinates as input, such as $\bm{\varphi}_{1}$, $\bm{\varphi}_{2}$ and $\bm{\theta}$, and they have difficulty learning high-frequency information [22].
-In CFNet, there are many MLP structures and transformer structures containing MLPs, so such a problem will degrade instance segmentation performance.
-To alleviate this problem, inspired by Gaussian positional encoding (GPE) for coordinates [20], CFNet introduces adaptive GPE (AGPE) to map the coordinate set ($\bm{\varphi}_{1},\bm{\varphi}_{2},\bm{\theta}$) into random Fourier features to better learn high-frequency information and enhance the performance of coordinate-based MLPs.
+In OBSeg, there are many MLP structures and transformer structures containing MLPs, so such a problem will degrade instance segmentation performance.
+To alleviate this problem, inspired by Gaussian positional encoding (GPE) for coordinates [22], OBSeg introduces adaptive GPE (AGPE) to map the coordinate set ($\bm{\varphi}_{1},\bm{\varphi}_{2},\bm{\theta}$) into random Fourier features to better learn high-frequency information and enhance the performance of coordinate-based MLPs.
 Given a d-dimensional coordinate $\bm{\tau}\in{[0,1)}^{d}$, the dense Fourier feature mapping is
@@
 To represent position information ($\gamma(\bm{\varphi}_{1})$, $\gamma(\bm{\varphi}_{2})$) and orientation information ($\gamma(\bm{\theta})$) in a unified representation space, we set $l_{p}=l_{\theta}$.
-Figure 3: Architecture of the proposed OBB prompt encoder.
-The input is an OBB ($x,y,w,h,\theta$), where $(x,y)$, $w$, $h$ and $\theta$ represent the center point, width, height and orientation, respectively.
 In many HBB prompt encoders of BSMs, $(\bm{\rho}{\bm{\varphi}_{1}},\bm{\rho}{\bm{\varphi}_{2}},\bm{\rho}{\bm{\theta}})$ and ($\bm{\bm{\Sigma}}_{\bm{\varphi}_{1}},\bm{\bm{\Sigma}}_{\bm{\varphi}_{2}},\bm{\bm{\Sigma}}_{\bm{\theta}}$) are considered as hyperparameters, which leads to difficulty in adapting to changes in datasets and searching optimal parameter settings.
-In CFNet, we adaptively learn ($\bm{\bm{\Sigma}}_{\bm{\varphi}_{1}},\bm{\bm{\Sigma}}_{\bm{\varphi}_{2}},\bm{\bm{\Sigma}}_{\bm{\theta}}$).
+In OBSeg, we adaptively learn ($\bm{\bm{\Sigma}}_{\bm{\varphi}_{1}},\bm{\bm{\Sigma}}_{\bm{\varphi}_{2}},\bm{\bm{\Sigma}}_{\bm{\theta}}$).
 $(\bm{\rho}{\bm{\varphi}_{1}},\bm{\rho}{\bm{\varphi}_{2}},\bm{\rho}{\bm{\theta}})$ are fixedly set to 0.
 Concretely, ($\bm{\bm{\Sigma}}_{\bm{\varphi}_{1}},\bm{\bm{\Sigma}}_{\bm{\varphi}_{2}},\bm{\bm{\Sigma}}_{\bm{\theta}}$) are the weights of a learnable embedding layer.
-Inspired by SAM, the mapped random Fourier features ($\gamma(\bm{\varphi}_{1})$ and $\gamma(\bm{\varphi}_{2})$) are summed with two learned embeddings ($\bm{\omega}_{1}$ and $\bm{\omega}_{2}$) that represent the top-left corner point and bottom-right corner point in CFNet, respectively.
+Inspired by SAM, the mapped random Fourier features ($\gamma(\bm{\varphi}_{1})$ and $\gamma(\bm{\varphi}_{2})$) are summed with two learned embeddings ($\bm{\omega}_{1}$ and $\bm{\omega}_{2}$) that represent the top-left corner point and bottom-right corner point in OBSeg, respectively.
 $\mathcal{E}(\bm{\varphi}_{1})=\gamma(\bm{\varphi}_{1})+\bm{\omega}_{1},$ (9)
@@
 $\mathcal{E}(\bm{\varphi}_{p})=\mathrm{Concat}[\mathcal{E}(\bm{\varphi}_{1}),~{}~{}\mathcal{E}(\bm{\varphi}_{2})].$ (11)
+Figure 3: Architecture of the proposed OBB prompt encoder.
+The input is an OBB ($x,y,w,h,\theta$), where $(x,y)$, $w$, $h$ and $\theta$ represent the center point, width, height and orientation, respectively.
 On the other hand, the parameterized orientation representation ($\sin\theta,\cos\theta$) of an OBB suffers from the boundary discontinuity problem [27].
 Due to the periodicity of orientations, at boundary orientations (such as $0^{\circ}$ and $180^{\circ}$), small changes in orientations will result in large jumps in ($\sin\theta,\cos\theta$).
-To alleviate the boundary discontinuity problem, CFNet introduces learnable orientation correction embeddings $\bm{\omega}_{\theta}$.
+To alleviate the boundary discontinuity problem, OBSeg introduces learnable orientation correction embeddings $\bm{\omega}_{\theta}$.
 At the orientations where the boundary discontinuity problem occurs, $\bm{\omega}_{\theta}$ adaptively adjusts the effect of this problem.
@@
 $\mathcal{P}_{\mathrm{OBB}}=\mathrm{Concat}[\mathcal{E}(\bm{\varphi}_{p}),~{}~{}\mathcal{E}(\bm{\theta})].$ (13)
-Figure 4: The process of knowledge distillation for the OBB prompt encoder and mask decoder.
-“TE”, “BE” and “OE” represent encoded feature embeddings with respect to the top-left point, bottom-right point and orientation of an OBB, respectively.
-“GS” stands for Gaussian smoothing.
 III-C Knowledge Distillation
 BSMs exhibit powerful performance, however they usually have high computational complexity.
-Some methods, such as MobileSAM [19], focus on distilling image encoders.
-However, the performance gap between the distilled model and the teacher model is still large.
-To tackle this, CFNet uses the distilled image encoders, and further distills the prompt encoders and mask decoders of BSMs.
-Considering the powerful data-fitting capability of the teacher model, the student model only imitates the teacher model and does not learn from GT labels.
+Although some lightweight BSMs, such as EfficientSAM, MobileSAM, and EdgeSAM, perform knowledge distillation on BSMs, the performance gap between the distilled model and the teacher model is still large.
+To tackle this, OBSeg introduces a Gaussian smoothing method to enhance distillation performance.
+Considering the powerful data-fitting capability of the teacher model, the student model only mimics the teacher model and does not learn from GT labels.
 Furthermore, to further enhance the performance and generalization capability of the student model, Gaussian smoothing is applied to teacher model outputs.
-The process of knowledge distillation for the proposed OBB prompt encoder and mask decoder is shown in Fig.
+While current methods focus on reducing the complexity of the image encoder, we focus on further improving the distillation performance of the OBB prompt encoder and mask decoder.
+We use a pretrained lightweight image encoder and perform distillation on the OBB prompt encoder and mask decoder, as shown in Fig.
 4.
-We define a one-dimensional Gaussian kernel $\mathbf{G}^{1}_{k}(x_{i};\sigma)$ with a length of $k$, which is given as
+We first define a one-dimensional Gaussian kernel $\mathbf{G}^{1}_{k}(x_{i};\sigma)$ with a length of $k$, which is given as
 $\mathbf{G}^{1}_{k}(x_{i};\sigma)=\frac{1}{\mathbf{G}^{1}_{s}}e^{-[\frac{(x_{i}-x_{c})^{2}}{2\sigma^{2}}]},$ (14)
@@
 […]
+Figure 4: The process of knowledge distillation for the OBB prompt encoder and mask decoder.
+“TE”, “BE” and “OE” represent encoded feature embeddings with respect to the top-left point, bottom-right point and orientation of an OBB, respectively.
+“GS” stands for Gaussian smoothing.
 Then, the corresponding encoded feature embeddings ($\mathcal{E}^{s}(\bm{\varphi}_{1}),\mathcal{E}^{s}(\bm{\varphi}_{2}),\mathcal{E}^{s}(\bm{\theta})$) of the student model mimic the smoothed feature embeddings ($\bm{G}^{t}(\bm{\varphi}_{1}),\bm{G}^{t}(\bm{\varphi}_{2}),\bm{G}^{t}(\bm{\theta})$) of the teacher model.
@@
 $+\frac{1}{2l}{||\bm{G}^{t}(\bm{\theta})-\mathcal{E}^{s}(\bm{\theta})||}^{2}_{2}.$ (19)
+Figure 5: Some visualization results of OBSeg’s predictions on the iSAID dataset. Segmentation masks of different colors represent different instances.
 For knowledge distillation on the mask decoder of a BSM, Gaussian smoothing makes the transition between boundary regions of different classes smoother and improves generalization capability.
@@
 where $\lambda$ is the trade-off factor between the distillation loss of the OBB prompt encoder and that of the mask decoder.
+TABLE I: Comparison results on the iSAID validation set. “RB Mask R-CNN” represents Rotated Blend Mask R-CNN (same elsewhere)
+Method Backbone AP AP50 AP75 APS APM APL FPS
+SOLO [28] Inception-ResNet 24.7 44.9 24.2 7.9 32.7 47.8 13.2
+Box2Mask-C [29] ResNet-101-FPN 26.6 50.6 23.8 10.6 33.6 47.4 11.1
+Luo et al. [30] ResNet-50-FPN 29.4 54.5 27.8 15.5 37.8 42.0 10.7
+OIS [7] ResNet-50-FPN 34.1 57.2 34.2 18.5 42.1 49.9 11.0
+PointRend [31] ResNet-50-FPN 35.6 59.0 37.3 20.3 44.8 52.9 15.7
+Mask Scoring R-CNN [32] ResNet-50-FPN 35.9 57.7 38.4 20.8 44.3 51.5 9.8
+DCTC [33] DLASeg 36.3 58.2 38.0 21.2 45.2 47.3 23.9
+SCNet [34] ResNet-50-FPN 37.3 59.5 40.3 23.3 44.8 52.3 7.6
+Hybrid Task Cascade [35] ResNet-50-FPN 37.4 60.2 40.1 23.5 44.6 53.5 5.8
+Ye et al. [36] ResNet-50-FPN 37.6 60.8 41.4 22.4 47.7 53.4 5.3
+RB Mask R-CNN [6] ResNet-50-FPN 37.9 60.6 40.9 23.0 46.8 52.7 6.1
+R-ARE-Net [37] DarkNet-19 38.0 60.8 40.7 25.2 45.9 52.7 -
+RB Mask R-CNN [6] ViT-Tiny 38.5 61.5 41.4 24.0 47.3 53.3 3.8
+ISOP [9] ResNet-101-FPN 39.4 64.2 41.7 24.9 48.6 55.4 16.6
+Ye et al. [36] Swin-Tiny 40.2 65.1 44.3 25.3 49.9 57.3 2.9
+ISOP [9] ViT-Tiny 40.3 65.2 43.9 25.6 49.4 56.8 9.1
+Finetuned MobileSAM [17] ViT-Tiny 33.4 58.8 32.9 19.1 42.5 47.8 21.5
+Finetuned EfficientSAM [16] ViT-Tiny 34.8 60.0 34.0 20.3 43.7 48.6 21.5
+Finetuned SAM [10] ViT-H 37.2 62.3 36.7 22.9 45.9 51.1 2.8
+RSPrompter-anchor [11] ViT-H 40.9 64.0 44.2 24.3 49.7 57.0 2.7
+RSPrompter-query [11] ViT-H 41.6 64.4 44.6 24.8 50.2 57.6 2.7
+OBSeg* ViT-H 45.1 69.4 48.3 28.8 55.1 60.7 2.7
+[…]
+OBSeg ViT-Tiny 43.7 65.2 46.1 25.7 51.4 59.6 20.4
 IV Experiments
+To verify the effectiveness of OBSeg, we conducted experiments on multiple public datasets.
+Since some datasets do not provide the GT parameters of OBBs, the smallest OBBs of instance masks are used as regression targets.
 IV-A Dataset
-To verify the effectiveness of the proposed methods, we conduct completely occluded object instance segmentation and dense object instance segmentation experiments on self-collected industrial and public datasets, respectively.
-Figure 5: Self-designed robotic system for completely occluded object instance segmentation in the industrial robot assembly environment of the large commercial aircraft C919.
-IV-A1 Industrial Dataset for Completely Occluded Objects
-The self-designed industrial robotic vision system for large commercial aircraft assembly is shown in Fig.
-5.
-It mainly consists of a UR10 robot, a vision sensor and a computer.
-Specifically, the UR10 robot is used to carry the vision sensor for operations.
-The computer is responsible for receiving and processing the image data collected by the vision sensor, and it runs the algorithm for completely occluded object instance segmentation.
-Due to the limited working space, reference holes are usually observed from the side.
-All the images in the dataset are collected by this system, including three types of occluded reference holes, i.e., reference holes with bolts, nuts and untreated bolts (see the left image of Fig.
-1).
-To reflect complex and changeable environments in the assembly process, each type of reference hole is photographed with the camera at various perspectives, distances and lighting conditions.
-We obtain a total of 800 images, with approximately the same number of each type.
-These images are divided into a training set, a validation set, and a testing set in a ratio of $7\colon 1\colon 2$ in a uniform sampling manner.
-All images are cropped to $1024\times 1024$.
-Then, these images are first flipped horizontally and vertically to quadruple the number.
-Experimental GT labels are obtained from industrial standard workpieces.
-Furthermore, the Mosaic method that randomly mixes four images is used for data augmentation.
-IV-A2 Public Dataset for Dense Objects
+IV-A1 iSAID
 iSAID [38] is a large-scale and densely annotated instance segmentation dataset in aerial images.
-It has 655,451 object instances across 15 categories.
-The training set, validation set and testing set contain 1411, 458 and 937 images, respectively.
-All images are cropped into $1024\times 1024$ patches with a stride of 500.
-Since iSAID dataset does not provide the GT parameters of OBBs, the smallest OBBs of instance masks are used as regression targets.
-We only choose densely distributed large vehicles, small vehicles and ships as targets for dense object instance segmentation.
-Figure 6: Some visualization results on the completely occluded object instance segmentation dataset.
-Bounding boxes of different colors and segmentation masks of different colors represent different object classes and different instances, respectively.
+It contains 2806 high-resolution aerial images and 655,451 object instances across 15 categories.
+The categories include airplanes, ships, storage tanks, baseball fields, tennis courts, basketball courts, ground track and field, ports, bridges, large vehicles, small vehicles, helicopters, roundabouts, swimming pools, and football fields.
+The training set, validation set and test set contain 1411, 458 and 937 images, respectively.
+All images are cropped into $1024\times 1024$ with a stride of 500.
+IV-A2 NWPU VHR-10
+NWPU VHR-10 instance segmentation dataset [39] is a widely-used dataset in geospatial tasks, which contains 800 high-resolution color and panchromatic infrared images.
+It has 650 very-high-resolution (VHR) optical images with objects and 150 VHR optical images with pure background.
+We randomly divide the dataset into the training set and the test set in a ratio of $7\colon 3$.
+IV-A3 PSeg-SSDD
+PSeg-SSDD dataset [40] supports ship instance segmentation and ship detection tasks.
+It contains 1160 SAR images with a resolution ranging from 1 to 15m, and there are various polarizations, resolutions, and scenes.
+We randomly divide this dataset into a training set of 928 images and a test set of 232 images.
+Figure 6: Some comparison results on the iSAID dataset. Segmentation masks of different colors represent different instances.
 IV-B Implementation Details
-The training settings for Oriented R-CNN [14] are the same as the original method.
+In this work, OBSeg chooses SAM as the segmentation foundation model.
+For OBB detection, we directly employ Oriented R-CNN [12], which is a two-stage OBB detection model with competitive detection performance.
+The training settings for Oriented R-CNN with ResNet-34 [41] are the same as the original method.
 In the OBB prompt encoder, the length of the sampled Fourier basis frequencies is 128.
-For knowledge distillation, the teacher model (called CFNet*) is based on SAM with the powerful ViT (vision transformer)-H image encoder and the proposed OBB prompt encoder.
-The student model (i.e., CFNet) uses a lightweight pretrained ViT-Tiny image encoder and the proposed OBB prompt encoder.
+For knowledge distillation, the teacher model (called OBSeg*) is based on SAM with the powerful ViT-H (vision transformer [42]) image encoder and the proposed OBB prompt encoder.
+The student model (i.e., OBSeg) directly uses the pretrained lightweight ViT-Tiny image encoder in MobileSAM and the proposed OBB prompt encoder.
 Their mask decoders and OBB detection modules are the same.
 During training, the weights of their image encoders are frozen, and the other parts are fine-tuned.
+Fine-tuning and distillation are performed for 5 and 8 epochs, respectively.
 The teacher model has the same training loss as SAM and has been trained before distillation.
@@
 More implementation details can be found in our open-source code.
-All experiments are conducted on an RTX 3080Ti GPU.
-For evaluation, we use the standard COCO metrics: AP (i.e., AP50-95), AP50 and AP75, where “AP” means average precision over IoU (Intersection over Union) threshold.
-IV-C Completely Occluded Object Instance Segmentation
-In completely occluded object instance segmentation experiments, we compare CFNet, SAM, and amodal methods.
-SAM uses the ViT-H to encode images and the HBB detection branch of Mask R-CNN to generate HBB prompts (the same below).
-Amodal methods directly predict occludee instances.
-CFNet and SAM perform the proposed instance segmentation method on occluders to obtain occludee instances.
-We also include SAM after fine-tuning on this dataset and the teacher model (CFNet*) used for distillation in comparison.
-For fair comparison, we replace the backbones of current amodal instance segmentation methods with the ViT-H image encoder used in BSMs (indicated by $\dagger$).
-TABLE I: Comparison results on the completely occluded object instance segmentation dataset.
-Method AP AP50 AP75 FPS
-AmodalMask [15] 19.49 35.18 24.36 16.1
-ASN [5] 40.88 61.24 45.93 18.6
-ORCNN [6] 37.60 57.78 40.16 22.2
-BCNet [4] 45.30 65.13 51.79 7.4
+All experiments are conducted on an RTX 3090 GPU.
+For evaluation, we use the standard evaluation metrics: AP (i.e., AP50-95), AP50, AP75, APS, APM and APL, where “AP” means average precision over IoU (Intersection over Union) threshold.
+TABLE II: Comparison results on the NWPU VHR-10 test set
+Method Backbone AP AP50 AP75 APS APM APL
+YOLACT [43] ResNet-50-FPN 43.3 77.5 44.0 23.1 40.5 54.3
+Mask Scoring R-CNN [32] ResNet-50-FPN 58.4 92.1 63.2 46.7 56.9 63.5
+PointRend [31] ResNet-50-FPN 61.1 90.1 64.7 52.8 59.9 61.0
+PointRend [31] ViT-Tiny 61.9 91.2 65.8 53.4 61.7 61.6
+Hybrid Task Cascade [35] ResNet-50-FPN 62.8 92.9 69.1 49.8 63.5 65.8
+ARE-Net [37] ResNet-101-FPN 64.8 93.2 71.5 53.9 65.3 72.9
+Shi and Zhang [44] ResNet-FPN 65.2 94.9 72.1 49.4 65.7 71.2
+FB-ISNet [45] DLA-BiFPN 67.0 94.5 72.0 - - -
+DCTC [33] DLASeg 67.7 91.8 75.5 56.4 66.4 69.9
+Ye et al. [36] Swin-Tiny 68.2 - - 50.9 68.4 65.7
+Ye et al. [36] ViT-Tiny 68.2 - - 51.4 68.0 65.8
+Finetuned MobileSAM [17] ViT-Tiny 60.8 91.0 63.1 51.4 61.5 60.7
+Finetuned EfficientSAM [16] ViT-Tiny 62.6 92.5 65.9 51.8 63.0 63.1
+Finetuned SAM [10] ViT-H 65.0 95.0 70.4 53.6 65.2 73.0
+RSPrompter-anchor [11] ViT-H 66.1 92.7 70.6 - - -
+RSPrompter-query [11] ViT-H 67.5 91.7 74.8 - - -
+OBSeg* ViT-H 71.1 96.7 77.4 59.0 70.5 74.6
 […]
-SAM [13] 26.15 42.74 30.28 2.8
-Fine-tuned SAM 68.66 82.23 71.38 2.8
-CFNet* 75.13 89.21 77.98 2.7
-CFNet 73.46 86.48 76.20 20.4
-As shown in Table I, the accuracy of CFNet and CFNet* is superior to other methods, especially for high-precision metrics.
-Additionally, CFNet exhibits competitive inference speed.
-Some visualization results are presented in Fig.
-6.
-Compared with amodal instance segmentation methods, CFNet provides an effective solution for completely occluded object instance segmentation.
-In addition, compared with fine-tuned SAM using HBBs, the accuracy of CFNet and CFNet* using OBBs is significantly improved, which indicates the necessity of using OBBs.
-Moreover, our proposed knowledge distillation method significantly reduces the computational complexity with a minor loss in accuracy.
-Figure 7: Some visualization results on the dense object instance segmentation dataset iSAID.
-Bounding boxes of different colors and segmentation masks of different colors represent different object classes and different instances, respectively.
-IV-D Dense Object Instance Segmentation
-For dense object instance segmentation, we compare CFNet, SAM (with HBBs), and instance segmentation methods using OBBs.
-We also include fine-tuned SAM and the teacher model used for distillation in comparison, and replace the backbones of current instance segmentation methods using OBBs with the ViT-H image encoder used in BSMs.
-TABLE II: Comparison results on the dense object instance segmentation dataset. “RB Mask R-CNN” represents Rotated Blend Mask R-CNN [9].
-Method AP AP50 AP75 FPS
-OIS [10] 27.41 47.85 30.71 18.4
-RB Mask R-CNN [9] 29.07 50.96 32.64 16.3
-ISOP [12] 34.73 59.16 36.84 19.6
+OBSeg ViT-Tiny 69.5 95.1 75.6 57.7 69.2 73.3
+TABLE III: Comparison results on the PSeg-SSDD test set
+Method Backbone AP AP50 AP75 APS APM APL
+YOLACT [43] ResNet-50-FPN 57.8 91.4 70.9 58.4 56.5 58.6
+SA R-CNN [46] ResNet50-FPN 59.4 90.4 77.6 63.3 62.5 47.7
+EMIN [47] - 61.7 94.3 76.8 62.1 61.3 40.1
+FL-CSE-ROIE [48] ResNet-101-FPN 62.6 93.7 78.3 63.3 61.2 75.0
+MAI-SE-Net [49] ResNet-101-FPN 63.0 94.4 77.6 63.3 62.5 47.7
+LFG-Net [50] ResNeXt-64 64.2 95.0 81.1 63.1 68.2 43.1
+Mask Scoring R-CNN [32] ResNet-50-FPN 65.4 94.1 81.9 66.4 62.0 40.5
+PointRend [31] ResNet-50-FPN 65.6 94.5 82.3 67.0 62.2 16.8
+Hybrid Task Cascade [35] ResNet-50-FPN 66.1 95.4 83.2 66.2 68.9 29.4
+PointRend [31] ViT-Tiny 66.2 95.2 83.3 67.7 64.0 20.4
+DCTC [33] DLASeg 67.4 93.9 83.8 64.4 76.6 80.2
+Shi and Zhang [44] ResNet-FPN 67.8 97.5 85.0 67.7 69.1 58.9
+Ye et al. [36] Swin-Tiny 68.4 - - 69.8 63.7 41.0
+Ye et al. [36] ViT-Tiny 68.5 - - 69.4 64.6 42.5
+Finetuned MobileSAM [17] ViT-Tiny 62.9 93.6 77.1 64.0 64.8 49.9
+Finetuned EfficientSAM [16] ViT-Tiny 63.5 94.5 79.6 64.3 65.2 54.3
+Finetuned SAM [10] ViT-H 66.5 95.0 83.3 64.4 66.7 58.2
+RSPrompter-anchor [11] ViT-H 66.8 94.7 84.0 - - -
+RSPrompter-query [11] ViT-H 67.3 95.6 84.3 - - -
+OBSeg* ViT-H 71.8 97.0 87.5 71.2 75.3 72.4
 […]
-SAM [13] 14.37 41.88 18.52 2.8
-Fine-tuned SAM 20.26 45.30 24.59 2.8
-CFNet* 45.10 66.45 47.92 2.7
-CFNet 43.76 64.29 46.13 20.4
-As presented in Table II, CFNet and CFNet* outperform other methods, and CFNet achieves the highest inference speed of 20.4 FPS.
-In addition, CFNet achieves a remarkable dense object instance segmentation accuracy of 43.76% AP, 64.29% AP50, and 46.13% AP75.
+OBSeg ViT-Tiny 69.8 96.1 85.4 69.4 73.1 66.6
+IV-C Comparisons with the State-of-the-Art
+IV-C1 Comparison Experiments on the iSAID Dataset
+To evaluate the effectiveness of OBSeg, we compare OBSeg with current instance segmentation methods using OBBs on the iSAID dataset.
+We also include some instance segmentation methods using HBBs (e.g., SAM after finetuned on this dataset) and other SAM-based methods using HBBs (e.g., RSPrompter) for comparison.
+For fairness, SAM and the teacher model OBSeg* use ViT-H as the image encoder, and OBSeg uses ViT-Tiny as the image encoder.
+We also replace the backbones of some non-BSM-based methods with pretrained ViT-Tiny.
+SAM, MobileSAM, and EfficientSAM apply powerful YOLOv7 [51] as the HBB detector.
+$\dagger$” indicates that Gaussian smoothing is not applied during distillation, similar to its use in current lightweight BSMs.
+The inference is conducted on the RTX 3090 GPU, and the inference speed of some models is based on the results we reproduced.
+As presented in Table I, OBSeg* and OBSeg significantly outperform other methods, achieving 45.1% AP and 43.7% AP, respectively.
+Compared with instance segmentation methods using HBBs, methods using OBBs perform better since OBBs are more suitable for segmenting densely packed objects.
+Significant accuracy improvements over fine-tuned SAM series and other SAM-based methods using HBBs also indicate the benefit of using OBBs.
+In addition, OBSeg is less dependent on OBB detection performance and outperforms existing “segmentation within bounding box” methods by a large margin.
 Fig.
-7 shows some dense object instance segmentation results.
-Compared with existing instance segmentation methods using OBBs, CFNet achieves higher accuracy since it reduces the dependence on OBBs and is based on BSMs.
-Significant accuracy improvements over fine-tuned SAM also indicate the benefit of using OBBs.
-Moreover, compared with CFNet*, the notable inference speed improvement and minor accuracy loss of CFNet also demonstrate the effectiveness of the proposed knowledge distillation method.
-IV-E Ablation Studies
-To evaluate the effectiveness of the components of CFNet, a series of ablation experiments are conducted on the iSAID dataset. We use the most commonly used AP50 as the metric.
-IV-E1 Oriented Bounding Box Prompt Encoder
+6 shows some instance segmentation visualization results of OBSeg and current methods.
+For dense objects with oblique orientations, the segmentation results of methods using HBBs incorrectly include other instances, reducing the instance segmentation performance.
+Furthermore, when predicted OBBs significantly deviate from GT OBBs, the segmentation performance of existing “segmentation within bounding box” methods is substantially affected.
+Based on the OBB prompt mechanism, OBSeg still accurately segments instances when the OBB predictions are inaccurate.
+Compared with other BSMs, OBSeg is very lightweight.
+Even though SAM has a powerful ViT-H image encoder, its performance is still inferior to OBSeg.
+Among BSMs of the same lightweight level, OBSeg achieves higher accuracy and a competitive inference speed of 20.4 FPS, which is also better than most methods not based on BSMs.
+This demonstrates that OBSeg is more efficient than current instance segmentation methods.
+Moreover, compared with OBSeg*, the notable inference speed improvement and minor accuracy loss of OBSeg also demonstrate the effectiveness of the proposed knowledge distillation method.
+Gaussian smoothing further enhances the distillation performance and generalization capability.
+More instance segmentation results of OBSeg on the iSAID dataset are displayed in Fig.
+5.
+IV-C2 Comparison Experiments on the NWPU VHR-10 Dataset
+We further verify the instance segmentation performance of OBSeg on the NWPU VHR-10 dataset.
+For fairness, we compare instance segmentation methods using HBBs, methods using OBBs, and BSM-based methods.
+As in the setting on the iSAID dataset, the HBB detectors in the BSM series all use the powerful YOLOv7.
+We also replace the backbones of some non-BSM-based methods with ViT-Tiny.
+[…]
+IV-C3 Comparison Experiments on the PSeg-SSDD Dataset
+Similar to the experimental settings of the iSAID and NWPU VHR-10 datasets, we compare methods that use HBBs, OBBs, and those based on BSMs for instance segmentation.
+As presented in Table III, similar to the comparison experiment results on the iSAID and NWPU VHR-10 datasets, OBSeg* and OBSeg show their superior performance.
+Before distillation, the teacher model OBSeg* achieves an accuracy of 71.8% AP.
+After distillation, OBSeg greatly retains the model’s performance, achieving an instance segmentation accuracy of 69.8% AP, which is superior to current methods.
+IV-D Ablation Studies
+To evaluate the effectiveness of the components of OBSeg, a series of ablation experiments are conducted on the iSAID dataset. We use the most commonly used AP50 as the metric.
+IV-D1 Oriented Bounding Box Prompt Encoder
 The OBB prompt encoder introduces positional encoding and specific embeddings to encode OBBs.
-In Table III, “OCE”, “TCE” and “BCE” represent the specific orientation correction embeddings, top-left corner embeddings and bottom-right corner embeddings, respectively.
+In Table IV, “OCE”, “TCE” and “BCE” represent the specific orientation correction embeddings, top-left corner embeddings and bottom-right corner embeddings, respectively.
 Compared with GPE, AGPE encodes location information better by adaptively learning parameters of Gaussian distributions.
-Specific embeddings further improve instance segmentation performance.
-TABLE III: Performance of each component of the OBB prompt encoder.
-GPE AGPE TCE+BCE OCE AP50
-61.40
-62.27
-63.56
-64.29
-TABLE IV: Comparison of different learning targets and label smoothing methods in knowledge distillation.
-GT+ULS GT+GS T+GT T T+ULS T+GS- T+GS
-63.51 63.72 64.01 64.07 64.15 64.19 64.24
-IV-E2 Knowledge Distillation
-We first compare the performance of CFNet learning from the teacher model and GT labels, and test the impact of different label smoothing methods.
+Moreover, specific embeddings further improve instance segmentation performance by 0.97% AP50.
+Orientation correction embeddings alleviate the periodicity problem of orientation.
+IV-D2 Knowledge Distillation
+We compare the performance of OBSeg learning from the teacher model and GT labels, and test the impact of different label smoothing methods.
 “T” represents learning from the teacher model.
 The smoothing factor in ULS is 0.1.
-Gaussian smoothing (GS) has the same settings as SVLS, with a Gaussian kernel size of $3\times 3$ and a standard deviation of 1.0.
+For distillation on the mask decoder, Gaussian smoothing (GS) has the same settings as SVLS, with a Gaussian kernel size of $3\times 3$ and a standard deviation of 1.0.
+For distillation on the OBB prompt encoder, GS has a Gaussian kernel length of 3 and a standard deviation of 0.3.
 The weight factor of the teacher model and GT labels is 0.1.
 “GS-” represents distillation only for the mask decoder.
-The experimental results in Table IV demonstrate that learning from the teacher model has higher performance than learning from GT labels (learning only from GT labels means no knowledge distillation is performed), which is mainly thanks to the powerful data-fitting capabilities of BSMs.
+The experimental results in Table V demonstrate that learning from the teacher model has higher performance than learning from GT labels (learning only from GT labels means no knowledge distillation is performed), which is mainly thanks to the powerful data-fitting capabilities of BSMs.
 Furthermore, adding distillation for the OBB prompt encoder effectively enhances performance.
-Moreover, Gaussian smoothing further improves instance segmentation performance and is superior to ULS.
-In addition, we compare the impact of different parameter settings in Gaussian smoothing on instance segmentation performance in Table V.
-When $k=5$, $\sigma=0.3$, and $\delta=1.0$, the highest instance segmentation accuracy of 64.29% AP50 is achieved.
-TABLE V: Impact of parameter settings in Gaussian smoothing.
+Moreover, compared with distillation methods that are not based on Gaussian smoothing in existing lightweight distilled BSMs (e.g., EfficientSAM and MobileSAM), Gaussian smoothing in OBSeg further improves distillation performance and is superior to ULS.
+In addition, we compare the impact of different parameter settings in Gaussian smoothing on instance segmentation performance in Table VI.
+When $k=5$, $\sigma=0.3$, and $\delta=1.0$, the highest instance segmentation accuracy of 65.24% AP50 is achieved.
+TABLE IV: Performance of each component of the OBB prompt encoder
+GPE AGPE TCE+BCE OCE AP50
+64.00
+64.27
+64.86
+65.24
+TABLE V: Comparison of different learning targets and label smoothing methods in knowledge distillation
+GT+ULS GT+GS T+GT T T+ULS T+GS- T+GS
+64.23 64.44 64.73 64.79 64.87 64.91 65.16
+TABLE VI: Impact of parameter settings in Gaussian smoothing
 $\sigma=0.3$ $\sigma=0.5$
 $\delta=0.5$ $\delta=1.0$ $\delta=0.5$ $\delta=1.0$
-$k=3$ 64.24 64.24 64.20 64.17
-$k=5$ 64.22 64.29 64.18 64.21
-$k=7$ 64.19 64.25 64.15 64.14
-IV-E3 Oriented Bounding Box Detection
-To explore the impact of the OBB detection module on the instance segmentation performance, we compare Oriented R-CNN used in CFNet with some other OBB detection methods [26, 27].
-In addition, a HBB detection method (HBB detection branch of Mask R-CNN) is also added for comparison.
-As shown in Table VI, the effective use of OBBs significantly improves instance segmentation accuracy, and CFNet is not sensitive to specific OBB detection methods.
-TABLE VI: Performance comparison of OBB and HBB detectors.
-Mask Oriented Oriented S2ANet [27]
-R-CNN [16] R-CNN [14] Reppoints [26]
-42.95 64.29 64.37 64.06
-V Conclusions
-This paper proposes CFNet, a unified coarse-to-fine framework for both completely occluded and dense object instance segmentation in robot vision measurement.
-CFNet uses BSMs and introduces a novel OBB prompt encoder.
-To make CFNet more lightweight, a knowledge distillation method involving Gaussian smoothing for teacher model outputs is applied to the OBB prompt encoder and mask decoder.
-Experiments on both industrial and public datasets demonstrate that CFNet outperforms current instance segmentation methods.
-One limitation of this work is that CFNet relies on prior geometric properties to achieve completely occluded object instance segmentation.
-In the future, we plan to utilize more types of prior knowledge to improve the robustness.
-It is worth noting that CFNet not only uniformly handles the two difficulties of completely occluded and dense object instance segmentation, but also can be used for general instance segmentation tasks.
-Moreover, CFNet can also handle objects that are simultaneously in completely occluded and dense environments, such as instance segmentation of containers with UAVs.
-We also plan to explore more applications of CFNet in various complex robotic tasks.
+$k=3$ 65.16 65.16 65.12 65.09
+$k=5$ 65.14 65.24 65.10 65.13
+$k=7$ 65.11 65.17 65.07 65.06
+TABLE VII: Performance comparison of OBB and HBB detectors
+YOLOv7 [51] Oriented Oriented S2ANet [52]
+R-CNN [12] Reppoints [53]
+58.81 65.24 65.37 65.09
+IV-D3 Oriented Bounding Box Detection
+To explore the impact of the OBB detection module on the instance segmentation performance, we compare Oriented R-CNN used in OBSeg with some other OBB detection methods, such as Oriented Reppoints [53] and S2ANet [52].
+In addition, the HBB detection method YOLOv7 is also added for fair comparison.
+As shown in Table VII, the effective use of OBBs significantly improves instance segmentation accuracy, and OBSeg is not sensitive to specific OBB detection methods.
+V Discussion
+V-1 Comparison with Other Methods Using BSMs
+Currently, there are some works that use BSMs, such as SAM, for instance segmentation tasks.
+These methods mainly aim to leverage the powerful segmentation capabilities of BSMs, which are acquired after training on large-scale datasets.
+However, these methods rely on HBB prompt input, which introduces many interference regions when performing instance segmentation on objects densely packed in multiple orientations, thereby reducing the segmentation performance.
+Different from these methods, our proposed OBSeg uses more accurate OBBs as prompts and achieves better instance segmentation performance.
+Moreover, we introduce Gaussian smoothing-based knowledge distillation to make OBSeg more lightweight.
+Experimental results on iSAID, NWPU VHR-10, and PSeg-SSDD datasets demonstrate that OBSeg is superior to current BSM-based methods in terms of instance segmentation accuracy and has competitive inference speed.
+V-2 Analysis of Inference Accuracy and Speed
+Compared with methods using HBBs or those not based on BSMs, OBSeg uses BSMs with OBB prompts, which have fewer interference regions and are less dependent on bounding box detection performance.
+This enables OBSeg to achieve higher instance segmentation accuracy on multiple datasets.
+In addition, since BSMs have powerful segmentation capabilities but slow inference speed, we have applied model lightweighting.
+To maintain high segmentation performance, we propose a knowledge distillation method based on Gaussian smoothing.
+Compared with existing methods, the lightweight distilled OBSeg has competitive inference speed while maintaining high segmentation accuracy.
+V-3 Limitations and Future Work
+Although OBSeg achieves better instance segmentation performance than existing methods, it also has some limitations.
+One limitation of this work is that the instance segmentation performance of OBSeg is not robust to objects with very large aspect ratios or extreme scales.
+This is mainly because these types of objects are in the minority in the dataset, and the model has insufficient feature recognition capabilities for objects with these characteristics.
+To address this type of long-tail distribution problem, future work will build upon the BSM framework by adopting class-balanced sampling strategies, extreme-scale feature learning, and adaptive adjustment of loss functions.
+VI Conclusions
+This paper proposes OBSeg, an accurate and fast instance segmentation framework using BSMs with OBB prompts in remote sensing images.
+OBSeg introduces a novel OBB prompt encoder, and we perform knowledge distillation to make OBSeg more lightweight and practical.
+Specifically, a Gaussian smoothing-based distillation method is introduced to improve the instance segmentation performance of lightweight distilled BSMs.
+Thanks to OBB prompts, OBSeg is less dependent on bounding box detection performance and has fewer interference regions than instance segmentation methods using HBBs or HBB prompts.
+Experiments on multiple public datasets, such as iSAID and NWPU VHR-10 datasets, demonstrate that OBSeg outperforms current instance segmentation methods in terms of instance segmentation accuracy and has competitive inference speed.
+One limitation of this work is that the instance segmentation performance of OBSeg is not robust to objects with very large aspect ratios or extreme scales.
+Future work will focus on improvements in data, extreme-scale feature learning, and loss functions.